OpenStudy (anonymous):

subtract x^2+6x/x^2-2x-8 -x^2-12/x^2-2x-8

OpenStudy (anonymous):

\[\frac{x^2+6 x}{x^2-2 x-8}-\frac{x^2-12}{x^2-2 x-8}=\frac{6}{x-4} \]